### v1

# Helical randomization of magnetized Galactic plasmas: from magnetorotational disc dynamo to the Faraday rotation sky

Abstract
It is shown, using results of numerical simulations and galactic observations that the transition from deterministic chaos to hard turbulence in the Galactic magnetized plasmas (global and those generated in the internal accretion disk in the high-energy surrounding of a supermassive black hole at the Galactic center) happens through a randomization process. The notion of distributed chaos has been used to describe the randomization process. The randomization can be quantified with the main parameter of the distributed chaos which in turn can be related to magnetic helicity or its dissipation rate. The magnetic fields can impose their level of randomization on the electron density. Results of the numerical simulations of the Galactic dynamos: the inner disk’s ones (based on the magnetorotational instability) and global ones, are in good agreement with this approach, as well as with the results obtained using observations of the Faraday rotation sky.
1 Introduction
The observational estimates of the magnetic fields of the Galaxy turned out to be much larger than those predicted for the primordial magnetic fields. Therefore different mechanisms of amplification of the seed magnetic fields by the intensive motion of the electrically conducting Galactic plasmas were suggested in the literature. The intensive chaotic/turbulent motion of the Galactic plasmas can be produced, for instance, by the differential rotation of the Galaxy and by the supernova explosions. The special role of a supermassive black hole at the Galactic center in supporting the intensive chaotic/turbulent motion in the inner accretion disk’s plasma should be also emphasized.
Numerous models (numerical simulations) of the global Galactic magnetohydrodynamic dynamo were suggested (some of the recent models will be discussed below).
For the weakly compressible central disk, the physical processes at kinetic scales become important and demand fully kinetic dynamo simulations (some of such recent numerical simulations will be also discussed below).
The numerical models (simulations) not only allow a better understanding of the relevant physical processes but they become indispensable because of a principal difficulty of measurements (observations) of Galactic magnetic fields. Till now technically all observables related to the Galactic magnetic fields are entangled and mixed with other physical characteristics of the magnetized plasmas (electron density, for instance).
Another principal difficulty in this area is the absence of a relevant theory of the chaotic/turbulent processes. The scaling approach, widely used for data interpretation, demands a rather wide range of scales for validation of its applicability that is scarcely achievable in practice.
In the absence of such a theory, the notion of smoothness can be used to classify (quantify) chaotic/turbulent dynamical regimes. For this purpose, one can use spectral analyses because the stretched exponential spectra
$E(k)\propto\exp-(k/k_{\beta})^{\beta}$ (1)
are typical for smooth chaotic dynamics. Here $k$ is the wavenumber and $01\geq\beta>$ 0.
The particular value $\beta=1$ typically characterizes the deterministic chaos (see, for instance, Frisch and Morf 8, Brandstater and Swinney 5, Maggs and Morales 18, 19, 20, Khurshid et al 15 and references therein):
$E(k)\propto\exp(-k/k_{c}).$ (2)
For $1>\beta$ the smooth chaotic dynamics is not deterministic (this dynamics will be called ‘distributed chaos’ and the term will be clarified below). The term “soft turbulence” suggested in a paper (Wu et al. 34) could be also appropriate.
The value of the $\beta$ can be used as a measure of randomization: the further the value of $\beta$ is from the $\beta=1$ corresponding to deterministic chaos, the more significant the randomization. The small value of $\beta$ can be considered as a precursor of hard turbulence (Wu et al. 34). For non-smooth dynamics (the hard turbulence Wu et al. 34) the power-law (scaling) spectra are typical.
Figure 1: Magnetic energy spectrum at the saturated stage of a small-scale dynamo for a sufficiently large magnetic Reynolds number $Re_{m}=3000$ and small Reynolds and Mach numbers $Re=100$, $M=0.122$
Figure 1, for instance, shows the magnetic energy spectrum obtained in a recent direct numerical simulation of a small-scale magnetohydrodynamic dynamo adapted to galaxy plasmas (Brandenburg et al. 4) (the spectral data were taken from Fig. 4 of the paper by 4).
The authors of the paper considered a weakly compressible turbulence with an isothermal equation of state with the pressure $p=\rho c_{s}^{2}$, where $\rho$ is the density and $c_{s}$ is the constant speed of sound. The dynamics of the magnetic field ${\bf B}=\nabla\times{\bf A}$ (where ${\bf A}$ is the magnetic vector potential), the velocity ${\bf u}$, and the density $\rho$ was described by the system of equations:
$\frac{\partial{\bf A}}{\partial t}={\bf u}\times{\bf B}+\eta\nabla^{2}{\bf A},$ (3)
$\frac{D{\bf u}}{Dt}=-c_{s}^{2}\nabla\ln\rho+\frac{1}{\rho}\left[{\bf J}\times{\bf B}+\nabla\cdot(2\rho\nu{\bf S})\right]+{\bf f},$ (4)
$\frac{D\ln\rho}{Dt}=-\nabla\cdot{\bf u},$ (5)
where the current density ${\bf J}=\nabla\times{\bf B}/\mu_{0}$ ($\mu_{0}$ is the vacuum permeability), $\nu$ is viscosity, ${\bf S}$ is the rate-of-strain tensor, and ${\bf f}$ is a random nonhelical forcing function.
Figure 1 shows the magnetic energy spectrum at the saturated stage of the small-scale dynamo for a sufficiently large magnetic Reynolds number $Re_{m}=3000$ and a small Reynolds and Mach numbers $Re=100$, $M=0.122$ (magnetic Prandtl number $Pr_{m}=30$). In this case, one can expect deterministic chaos in the magnetic field generated by the dynamo. And indeed the dashed curve in Fig. 1 indicates the best fit by the exponential spectrum Eq. (2) typical for the deterministic chaos. The dotted arrow indicates position of $k_{c}/k_{\nu}$ where $k_{\nu}=(\varepsilon/{\nu}^{3})^{1/4}$ is the Kolmogorov or viscous dissipation scale.
In recent papers (Seta and Federrath 30, Seta et al. 32) results of analogous simulations were reported. But in this case, the Mach number $M=10$ was reached. Figure 2 shows the magnetic energy spectra obtained in these simulations. The spectral data were taken from Fig. C4 of the Ref. (Seta et al. 32). One can see that for $M=0.1$ the dashed curve indicates the best fit by the exponential spectrum Eq. (2) typical for the deterministic chaos (cf Fig. 1), whereas for $M=10$ the dashed curve indicates the best fit by a stretched exponential spectrum Eq. (1) with $\beta=1/2$, i.e. the distributed chaos.
Figure 2: Magnetic energy spectrum at the saturated stage of a small-scale dynamo for $M=0.1$ and $M=10$.
It will be shown in the next Section 2 that the value of $\beta=1/2$ corresponds to the distributed chaos dominated by the magnetic helicity and the results of a fully kinetic 3D numerical simulation of a magnetorotational dynamo in the inner galactic disc will be interpreted in this vein (for the case of zero net magnetic helicity the spontaneous breaking of local reflectional symmetry will be used to justify this approach). Then in Section 3 a magneto-inertial range of scales dominated by magnetic helicity will be introduced in the frames of the Kolmogorov-Iroshnikov phenomenology (which is characterized by different values of $\beta$, see also Bershadskii 3). Section 4 will compare the theoretical considerations to the results of numerical simulations of the global galactic magnetic field and observations of the Galaxy’s Faraday rotation sky.
2 Distributed chaos and magnetic helicity
2.1 Magnetic helicity
The ideal magnetohydrodynamics has three quadratic (fundamental) invariants: total energy, cross and magnetic helicity (Moffatt and Tsinober 25). The robustness of magnetic helicity conservation depends on the magnetic Reynolds number value $Re_{m}$. Most astrophysical systems (including the galaxies) have very large magnetic Reynolds numbers (see, for instance, a recent paper by Zenati and Vishniac 35 and references therein).
The average magnetic helicity density can be defined as
$h_{m}=\langle{\bf a}{\bf b}\rangle$ (6)
where ${\bf b}=[{\nabla\times\bf a}]$ is the fluctuating magnetic field, and delimited-⟨$\langle$..$\rangle$ … ⟩ denotes a spatial average. The fluctuating magnetic field ${\bf b}$ and magnetic potential ${\bf a}$ have zero means.
The invariance of $h_{m}$ is not valid in a uniform mean magnetic field ${\bf B_{0}}$. To deal with this problem a generalized average magnetic helicity density $\hat{h}$ was considered in the paper (Matthaeus and Goldstein 22)
$\hat{h}=h_{m}+2{\bf B_{0}}\cdot\langle{\bf A}\rangle$ (7)
where ${\bf B}={\bf B_{0}}+{\bf b}$, ${\bf A}={\bf A_{0}}+{\bf a}$. It was shown that in the ideal magnetohydrodynamics (Matthaeus and Goldstein 22) (see also Shebalin 33)
$\frac{d\hat{h}}{dt}=0$ (8)
That means invariance of the generalized average magnetic helicity density $\hat{h}$ in a uniform mean magnetic field.
2.2 Distributed chaos driven by magnetic helicity
One can see in the Introduction (Fig. 2) that a change of physical parameters (in this case the Mach number) can result in a transition from deterministic chaos to distributed one (randomization).
Let us understand the transition. The change of physical parameters results in fluctuations of the characteristic scale $k_{c}$ in equation (2). One can take this phenomenon into account with the help of an ensemble averaging
$E(k)\propto\int_{0}^{\infty}P(k_{c})\exp-(k/k_{c})dk_{c}$ (9)
where the random fluctuations of $k_{c}$ are described by a probability distribution $P(k_{c})$. This is the reason why the non-deterministic smooth chaotic dynamics has been called ‘distributed chaos’.
One can use the dimensional considerations and the scaling relationship between $k_{c}$ and the characteristic value of the magnetic field $B_{c}$
$B_{c}\propto|h_{m}|^{1/2}k_{c}^{1/2}$ (10)
to find $P(k_{c})$ for the magnetic field dynamics dominated by the magnetic helicity.
For this purpose let $B_{c}$ be half-normally distributed $P(B_{c})\propto\exp-(B_{c}^{2}/2\sigma^{2})$ (Monin and Yaglom 26). The half-normal distribution is a normal distribution with zero mean truncated to only have nonzero probability density for positive values of its argument. For a normal random variable $B$, for instance, the variable $B_{c}=|B|$ has a half-normal distribution (Jonson et al. 12).
From Eq. (10) we then obtain
$P(k_{c})\propto k_{c}^{-1/2}\exp-(k_{c}/4k_{\beta})$ (11)
i.e. the chi-squared distribution (here $k_{\beta}$ is a new constant).
Substituting Eq. (11) into Eq. (9) we obtain
$E(k)\propto\exp-(k/k_{\beta})^{1/2}$ (12)
The average magnetic helicity density $h_{m}$ was used in the estimate Eq. (10) as an invariant. Taking into account that there are no physical processes effectively transferring the magnetic helicity to the resistive scales one can consider it as an adiabatic invariant (in comparison with the effectively dissipating magnetic and total energy) also in a weakly dissipative case (see for instance, Zenati and Vishniac 35).
2.3 Spontaneous breaking of local reflectional symmetry
In a special case of global (net) reflectional symmetry, the global (net) magnetic helicity is equal to zero. In contrast, the point-wise magnetic helicity is not (because of the spontaneous breaking of the local reflectional symmetry). This is an inherent property of chaotic/turbulent flows. The emergence of the blobs with non-zero magnetic/kinetic helicity accompanies the spontaneous local symmetry breaking (Moffatt and Tsinober 25, Moffatt 23, 24, Kerr 14, Holm and Kerr 9, Levich and Tsinober 17). The magnetic surfaces are defined by the boundary conditions: ${\bf b_{n}}\cdot{\bf n}=0$, where ${\bf n}$ is a unit normal to the boundary of a magnetic blob. The sign-defined magnetic helicity of any magnetic blob is an adiabatic invariant (Moffatt and Tsinober 25) (see also above). The magnetic blobs can be numbered with $H_{j}^{\pm}$ denoting the helicity of the j-blob and (‘+’ or ‘-’) denoting the blob’s helicity sign:
[…]
Then we can consider the adiabatic invariant
${\rm I^{\pm}}=\lim_{V\rightarrow\infty}\frac{1}{V}\sum_{j}H_{j}^{\pm}$ (14)
The summation here is over the blobs with a certain sign only (‘+’ or ‘-’), and $V$ is the entire volume of the blobs participating in the summation.
The adiabatic invariant ${\rm I^{\pm}}$ can be used instead of $h_{m}$ in the estimate Eq. (10) for the case of the local reflectional symmetry breaking
$B_{c}\propto|{\rm I^{\pm}}|^{1/2}k_{c}^{1/2}$ (15)
and the stretched exponential spectrum Eq. (12) can be obtained for this case as well.
Figure 3: Magnetic energy spectra (vs. the perpendicular to the magnetic field wavenumber $k_{\perp}$) obtained in a recent 3D fully kinetic particle-in-cell simulation of a magnetorotational disc dynamo in collisionless pair plasma at a quasi-stationary non-linear stage.
If a non-zero mean magnetic field ${\bf B}_{0}$ is not negligible the average magnetic helicity density $h_{m}$ could be replaced in the estimate Eq. (10) by the generalized averaged magnetic helicity density $\hat{h}$ Eq. (7) . Since the generalized averaged magnetic helicity density $\hat{h}$ has the same dimensionality as $h_{m}$ the magnetic energy spectrum will be the same stretched exponential Eq. (12) (analogously can be considered the case of the local reflectional symmetry breaking).
The stretched exponential spectrum Eq. (12) already appeared in Fig. 2 at the saturated stage of a small-scale dynamo for the Mach number $M=10$.
Figure 3 shows the magnetic energy spectra (vs. the perpendicular to the magnetic field wavenumber $k_{\perp}$) obtained in a recent fully kinetic particle-in-cell simulation of a magnetorotational disc dynamo in collisionless pair plasma at a quasi-stationary non-linear stage. The spectral data were taken from Fig 2a of a recent paper (Bacchini et al. 1).
This numerical simulation was related to the magnetorotational disk dynamo (based on the magnetorotational instability) in the high-energy surrounding of a supermassive black hole at the Galactic center (see Introduction). The kinetic scales can play a significant role in the physics of the collisionless plasmas therefore the fully kinetic numerical simulations are very important (about conservation of magnetic helicity at kinetic scales see, for instance Schekochihin et al. 31, Polygiannakis and Moussas 29, Bershadskii 3).
The dashed curves indicate the best fit by the stretched exponential Eq. (12) both for the toroidal and poloidal components of the magnetic field. The vertical dotted line indicates the positions of the ion and electron gyroradii which coincide in this case.
3 Magneto-inertial range of scales
An inertial range of scales is expected in hydrodynamic turbulence for the high Reynolds numbers. In the inertial range the energy spectra are supposed to be depending on the energy dissipation rate $\varepsilon$ only (Monin and Yaglom 26).
A magneto-inertial range of scales has been introduced recently in plasma dynamics for the magnetohydrodynamic and kinetic scales (Bershadskii 3). In this range of scales, two parameters: the total energy dissipation rate $\varepsilon$ (see above) and the magnetic helicity dissipation rate $\varepsilon_{h}$ (or dissipation rate of its modification $I^{\pm}$ in the case of global reflectional symmetry) govern the magnetic field dynamics.
There is an analogy of this approach to the Corrsin-Obukhov approach to the passive scalar in the turbulent flows where two governing parameters: the energy dissipation rate and the passive scalar dissipation rate, determined the so-called inertial-convective range (Monin and Yaglom 26) (see also Bershadskii and Sreenivasan 2).
According to this analogy, we can replace the estimate Eq. (10) with the estimate
$B_{c}\propto\varepsilon_{h}^{1/2}~{}\varepsilon^{-1/6}~{}k_{c}^{1/6}$ (16)
for the magneto-inertial range.
The estimates Eq. (10), (15), and (16) can be generalized
$B_{c}\propto k_{c}^{\alpha}$ (17)
At large $k_{c}$ the stretched exponential representation of the distributed chaos spectra
$\int_{0}^{\infty}P(k_{c})\exp-(k/k_{c})dk_{c}\propto\exp-(k/k_{\beta})^{\beta}$ (18)
can be employed for finding the probability distribution $P(k_{c})$ (Johnston 13)
$P(k_{c})\propto k_{c}^{-1+\beta/[2(1-\beta)]}~{}\exp(-\gamma k_{c}^{\beta/(1-\beta)})$ (19)
Using this asymptote a relationship between the parameters $\alpha$ and $\beta$ can be derived (using some simple algebra) from the Eqs. (17) and (19) in the case of the half-normally distributed $B_{c}$
$\beta=\frac{2\alpha}{1+2\alpha}$ (20)
Then for $\alpha=1/6$ (see Eq. (16)), one can find from Eq. (20) that
$E(k)\propto\exp-(k/k_{\beta})^{1/4},$ (21)
Figure 4: Magnetic energy spectra obtained in a 3D kinetic numerical simulation of magnetorotational dynamo based on the magnetorotational instability in a collisionless accretion disk at the saturated stage.
for the magneto-inertial range of scales.
Figure 4 shows the magnetic energy spectra obtained in a 3D kinetic numerical simulation of magnetorotational dynamo based on the magnetorotational instability in a collisionless accretion disk (cf Fig. 3). The spectral data were taken from Fig. 5a of a paper (Kunz et al. 16). The wavenumber $k$ was normalized by the disk height $H$. The magnetic Reynolds number was rather large $Re_{m}=37500$, ${\bf B}_{v}$ and ${\bf B}_{h}$ are vertical and horizontal magnetic field components, respectively.
The dashed curves indicate the spectral law Eq. (12) for ${\bf B}_{h}$ component and Eq. (21) for ${\bf B}_{v}$ component (an anisotropy of the level of randomization). The vertical dotted line indicates the position of the ion gyroradius $\rho_{i}$.
Figure 5 shows the power spectrum of the plasma density fluctuations at the saturated stage. The spectral data were taken from Fig 5b of the paper (Kunz et al. 16). The dashed curve indicates the spectral law Eq. (12). One can see that the horizontal component of the magnetic field imposes its level of the randomization $\beta=1/2$ on the plasma density fluctuations in this case (see also next Section).
4 Global galactic magnetic field and Faraday rotation sky
4.1 Numerical simulations
Figure 5: As in Fig. 4 but for density fluctuations.
Let us start with numerical simulations. In the Introduction, we have already discussed two simplified numerical simulations of the Galactic magnetic field (see Figs. 1 and 2). In a recent paper (Ntormousi et al. 27) generation (amplification) of the global magnetic field of a Milky-Way-like galaxy by the $\alpha-\omega$ dynamo was numerically simulated using a system of equations:
$\frac{\partial\rho}{\partial t}+\nabla(\rho{\bf u})$ $=$ 00 0 (22)
$\frac{\partial\rho{\bf u}}{\partial t}+\nabla\cdot\left(\rho\bf{u}\bf{u}-\bf{B}\bf{B}\right)+\nabla P_{tot}$ $=$ $-\rho\nabla\phi$ (23)
$\frac{\partial E_{tot}}{\partial t}+\nabla\left[~{}(E_{tot}+P_{tot})\bf{u}-(\bf{u}\cdot\bf{B})\cdot\bf{B}\right]$ $=$ $-{\bf u}\cdot\nabla\phi-\rho\Lambda+\Gamma$ (24)
$\frac{\partial\bf{B}}{\partial t}-\nabla\times(\bf{u}\times\bf{B})$ $=$ 00 0 (25)
$\nabla\cdot\bf{B}$ $=$ 00 0 (26)
where $\phi$ is the gravitational potential, $\Gamma=\Gamma(\rho,T)$ and $\Lambda=\Lambda(\rho,T)$ (as functions of density and temperature) are the heating and cooling rates of the plasma, the total pressure:
$P_{tot}=p+\frac{\bf{B}\cdot\bf{B}}{2}$ (27)
and the total energy
$E_{tot}=E_{int}+\rho\frac{\bf{u}\cdot\bf{u}}{2}++\frac{\bf{B}\cdot\bf{B}}{2}$ (28)
where the internal energy of the fluid is $E_{int}$. The equation of state for the plasma was taken as $P=(\gamma-1)E_{int}$.
Figure 6: Power spectrum of the chaotic/turbulent galactic magnetic field fluctuations (numerical simulations).
The model also included stars, stellar feedback, and a dark matter halo. The configuration of hydrodynamical fluid, dark matter, and stellar particles was chosen to simulate a Milky-Way-like galaxy. A weak toroidal magnetic field was taken as a seed initial condition.
In the simulation, turbulence is generated by the differential rotation of the galaxy and by the supernova explosions (supernova explosions produce a considerable amount of energy which then is transferred to the surrounding interstellar medium).
At a certain stage of the dynamo evolution, the magnetic field was smoothed at a given scale to separate the mean field and chaotic/turbulent residual fluctuations.
Figure 6 shows the magnetic energy spectrum of the chaotic/turbulent magnetic field fluctuations. The dashed curve indicates the best fit by the stretched exponential Eq. (21).
4.2 Observations
Now let us turn to observational results related to the Galactic magnetic field. The Faraday effect represents a rotation of the polarization position angle which propagates through magnetized plasma and entangles information on the plasma electron density $n_{e}$ with the line-of-sight (LoS) component of the magnetic field $B_{\mathrm{\parallel}}$ into the observable Faraday depth:
[…]
In a recent paper (Hutschenreuter et al. 10) results of disentangling an all-sky map of the average Galactic $B_{\mathrm{\parallel}}$ from the Faraday effect were reported. Several additional tracers of the $n_{e}$ (a $H-\alpha$ map Finkbeiner 7, pulsar data Manchester 21, the free-free map of the Planck survey Planck Collaboration 28, and extra-Galactic Faraday data Van Eak et al. 6) were used for this purpose and an all-sky map of the electron dispersion measure (the integrated electron density) was also constructed.
Figure 7: Power spectrum of the average $B_{\mathrm{\parallel}}$ sky map (inferred from the Farady rotation sky).
Figure 8: Power spectrum of electron dispersion measure sky map (inferred from the Farady rotation sky).
Figures 7 and 8 show the power spectra of the average $B_{\mathrm{\parallel}}$ and electron dispersion measure respectively. The spectral data were taken from Figs. 9a and 9b of the paper (Hutschenreuter et al. 10). In the original figures presented in the paper (Hutschenreuter et al. 10) the spectra are shown vs the spherical harmonic degree $l$. The spherical harmonic degree $l$ can be related to the wave number $k=\sqrt{l(l+1)}/R$, where $R$ is an effective radius of the Galactic map.
The dashed curves in Figs. 7 and 8 indicate the best fit by the stretched exponential spectral law Eq. (21). One can see that in this case the magnetic field imposes its degree of randomization on the electron dispersion measure.
5 Conclusions
The above-considered results obtained in numerical simulations (models) and inferred from the observations of the Faraday rotation sky show that magnetic helicity dominates the chaotic/turbulent dynamics (directly or through the Kolmogorov-Iroshnikov phenomenology) in the internal accretion disk plasma around the central black hole and the global Galactic plasma’s magnetic field and electron density. The magnetic field imposes its level of randomization on the plasma’s electron density .
This helcal randomization takes place even in the case of zero (or negligible) net helicity due to the phenomenon of spontaneous breaking of the local reflectional symmetry build-in the chaotic/turbulent plasma’s dynamics.
The applicability of the above-suggested approach to the kinetic scales (that is relevant to the internal disk’s collisionless plasma) has been also briefly discussed
Despite very considerable differences in the physical parameters and scales the results of the numerical simulations (models) related to the randomization phenomenon are in quantitative agreement with the observations of the Galactic plasmas.
6 Acknowledgments
I thank H.K. Moffatt, A. Pikovsky, and J.V. Shebalin for stimulating discussions.